OpenStudy (anonymous):

the linearization of f(x) = 5x^2 + 3x -2 near x =1 is (show work please)

OpenStudy (jamesj):

In general, given a function f which is differentiable at a point p, the linearization L of f at p is given by L(x-p) = f(p) + f'(p)(x-p) So using that formula, find L for your f and p.

OpenStudy (jamesj):

Actually for you, you'll want to use L(x) = f(p) + f'(p)(x-p) ----- (*) ===== The motivation for this formula is hopefully clear: the tangent line at p is the best linear approximation odor the graph of f. This equation (*) for a straight line is a the equation of that tangent line to the graph of f at that point.

OpenStudy (anonymous):

could you please explain the steps to solve, a little confused as to what variable represents what.

OpenStudy (jamesj):

For example, let's linearize f(x) = x^2 at x = 1. First f'(x) = 2x thus f'(1) = 2. Also f(1) = 1. Hence the linearization is L(x) = f(1) + f'(1)(x-1) = 1 + 2(x-1) = 2x - 1 This is the equation of the tangent line of y = x^2 at x = 1. For x = 1, L(1) = 1 = f(1), which is exactly what we'd expect.

OpenStudy (jamesj):

I.e., the linearization of f(x) = x^2 at p = 1 (or x = 1) is L(x) = f(p) + f'(p)(x-p) = 2x - 1

OpenStudy (anonymous):

ok got it trying it out thanks

OpenStudy (anonymous):

Answer : y = 13x - 7

OpenStudy (jamesj):

Let's see. If f(x) = 5x^2 + 3x -2, p = 1 then f'(x) = 10x + 3. f(p=1) = 6, f'(1) = 13 hence L(x) = 6 + 13(x-1) = 13x - 7 Yep.

OpenStudy (anonymous):

thanks